HELP ME PLEASE THANK YOU!!!!!

  1. Light coming from slit S1 makes distance r1to the screen. Light coming from slit S2 makes distance r2 to the screen. If the wave length is l, for the constructive interference r1 – r2have to be equal to

Bright fringe S o S,

  1. 3l;
  2. l;
  3. 3l/2;
  4. l/2;
  5. A and B;
  6. C and D.
Answer #1

For constructive interfernece, the path difference ( r1 - r2) is an even multiple of (0.5* wavelength) i.e. even multiple of (l / 2) .

( r1 - r2) = (2m) * ( l /2) = ( m* l) where m = 0, 1, 2, 3, .......

For constructive interference, ( r1 - r2) = l ;  ( r1 - r2) = 3l .

Option (A) and Option (B) both are correct.

Correct choice: (E)